OpenStudy (anonymous):

tiblada Group Title Sherri has a gross monthly salary of $4,700. She has a $425 car loan payment and a $330 student loan payment due each month. She also has two credit card loan payments each month, one of $127 and the other of $283. A bank will only give a mortgage loan if the total amount of monthly payments for all loans, including the mortgage, is no more than 36% of the monthly gross income. What is the maximum monthly mortgage loan payment for which Sherri can be approved?

OpenStudy (cj49):

36% of 4700 is 1692 current loan payments= 425+330+127+283=865 1692-865=827